OpenStudy (anonymous):

Solve the given system, using the substitution method. y = 3x – 7 6x – 2y = 12 There is no solution. (12, 14) (14, 12) There are an infinite number of solutions

OpenStudy (anonymous):

The eqn has no solution. \[a _{1} / b _{1} = a _{2} / b _{2} \neq a _{3} / b _{3}\] is the condition for no solution. a1 , a2,b1,b2 are coefficients before x and y respectively and a3, b3, are constants for the two eqns.

OpenStudy (anonymous):

Solve, using the substitution method. 3x + 2y = 11 x – 2 = – 4y

OpenStudy (anonymous):

Okay. \[x-2/ -4 = y\] \[2-x / 4 = y\] \[3x + 2 { ( 2-x) /4 } = 11\] \[3x + (2-x)/2 = 11\] \[6x + 2 -x = 22\] \[5x = 20\] \[x =4\]

OpenStudy (anonymous):

Solve, using the substitution method. y + 2x = 7 14 – 4x = 2y The solution is (1, 5) The solution is (21, 0) There are an infinite number of solutions. There is no solution.